OpenStudy (anonymous):

solve? 4^3 + 19 * 23-(17+20)

OpenStudy (anonymous):

464

OpenStudy (anonymous):

How did you get that answer? C:

OpenStudy (anonymous):

Start with the Parenthesis. 17 + 20 = 37. 4^3 + 19 * 23 - 37 Now, solve the exponent. 4 * 4 * 4 = 64 64 + 19 * 23 - 37 Now, solve multiplication. 19 * 23 = 437 64 + 437 - 37 Addition 64 + 437 = 501 Subtraction 501 - 37 = 464. :)

OpenStudy (anonymous):

OMG thank you!! :D

OpenStudy (anonymous):

You Welcome! =D

OpenStudy (anonymous):

:DD